# Improving Video Generation with Human Feedback

### v1

@@ Abstract
Video generation has achieved significant advances through rectified flow techniques, but issues like unsmooth motion and misalignment between videos and prompts persist. In this work, we develop a systematic pipeline that harnesses human feedback to mitigate these problems and refine the video generation model. Specifically, we begin by constructing a large-scale human preference dataset focused on modern video generation models, incorporating pairwise annotations across multi-dimensions. We then introduce VideoReward, a multi-dimensional video reward model, and examine how annotations and various design choices impact its rewarding efficacy. From a unified reinforcement learning perspective aimed at maximizing reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models. These include two training-time strategies: direct preference optimization for flow (Flow-DPO) and reward weighted regression for flow (Flow-RWR), and an inference-time technique, Flow-NRG, which applies reward guidance directly to noisy videos. Experimental results indicate that VideoReward significantly outperforms existing reward models, and Flow-DPO demonstrates superior performance compared to both Flow-RWR and standard supervised fine-tuning methods. Additionally, Flow-NRG lets users assign custom weights to multiple objectives during inference, meeting personalized video quality needs. Project page: https://gongyeliu.github.io/videoalign.
Machine Learning, ICML
[…]
@@ 1 Introduction
Recent advances in video generation have yielded powerful models (Kong et al., 2024; Polyak et al., 2024; Kuaishou, 2024; Brooks et al., 2024), capable of producing videos with convincing details and relatively coherent motion. Despite these notable achievements, current video generation systems still face persistent challenges, including unstable motion, imperfect alignment with user prompts, unsatisfactory visual quality, and insufficient adherence to human preferences (Zeng et al., 2024). In language modeling and image generation, learning from human preference (Ouyang et al., 2022; Zhou et al., 2023; Wallace et al., 2024) has proved highly effective in improving the response quality and aligning generative models with user expectations.
However, applying such preference-driven alignment strategies to video generation remains in its infancy. One key obstacle is the lack of large-scale, high-quality preference data. A pioneering work (He et al., 2024) introduces a human-rated video preference dataset, and concurrent studies (Wang et al., 2024b; Liu et al., 2024b; Xu et al., 2024a) have contributed additional annotations. Yet, these datasets primarily focus on videos generated by earlier open-source models, which are often of relatively low quality, contain artifacts, or remain limited in duration. With video generation techniques rapidly advancing, the gap between such datasets and the capabilities of state-of-the-art video diffusion models (VDMs) has become increasingly pronounced.
@@
A second challenge arises from the internal mechanisms of cutting-edge video generation models. Many modern systems employ rectified flow (Liu et al., 2022; Lipman et al., 2022), predicting velocity rather than noise. Recent studies (Wang et al., 2024b; Zhang et al., 2024a; Liu et al., 2024b; Xu et al., 2024a) have tested DPO (Rafailov et al., 2024; Wallace et al., 2024) and RWR (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024) on diffusion-based video generation approaches. However, adapting existing alignment methods to these flow-based models introduces new questions. While concurrent work (Domingo-Enrich et al., 2024) extended Diffusion-DPO to flow-based image generation, the aligned flow-based models in these explorations underperformed their unaligned counterparts.
To address these limitations, we present a comprehensive investigation into aligning advanced flow-based video generation models with human preferences. First, we select 12 advanced video generation models and curate a new large-scale multi-dimensional preference dataset of approximately 182k annotated examples. The dataset encompasses three critical dimensions—Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA)—to better capture how users evaluate generated videos. We then developed a multi-dimensional video reward model, systematically examining how annotations and different design choices affect its performance. Subsequently, we aggregated these multi-dimensional rewards into a comprehensive metric to capture overall human preferences. Furthermore, from a unified reinforcement learning perspective that maximizes reward with KL regularization, we introduce three alignment algorithms for flow-based models by extending those from diffusion models. These include two training-time strategies—Direct Preference Optimization for Flow (Flow-DPO) and Reward Weighted Regression for Flow (Flow-RWR)—and an inference-time technique, Flow-NRG. Our derivation reveals that the parameter $\beta$ (which modulates the KL divergence constraint) in Flow-DPO is dependent on the time step, resulting in Flow-DPO performing poorly on certain tasks. However, we found that by simply fixing $\beta$, Flow-DPO becomes the most effective alignment approach in our evaluations. Flow-NRG is an efficient inference-time alignment algorithm that applies reward guidance directly to noisy videos during the denoising process.It allows users to apply arbitrary weightings to multiple alignment objectives during inference, eliminating the need for extensive retraining to meet personalized user requirements.
@@
Figure 1: Overview of Our Video Alignment Paradigm. (a) Human Preference Annotation (Sec. 3.1). We construct a dataset of 182k (prompt, video A, video B) triplets, collecting preference annotations on Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA) from human evaluators. (b) Reward Mode Training (Sec. 3.2). We train a VLM-based reward model using the Bradley-Terry-Model-with-Ties formulation. (c) Video Alignment (Sec. 4). We adapt alignment techniques — DPO, RWR, and reward guidance — to flow-based video generation models and provide a comprehensive comparison of their effectiveness.
@@
  • Large-Scale Preference Dataset: We create a new 182k-sized human-labeled video generation preference dataset from 12 video generation models, capturing VQ, MQ, and TA dimensions.
@@
  • VideoGen-RewardBench: To more accurately assess video reward models, we have annotated prompt-video pairs from VideoGen-Eval. This dataset includes a wide variety of prompts and videos generated by modern VDMs. The resulting benchmark consists of 26.5k video pairs, each accompanied by a preference label.
@@
  • Flow-Based Alignment: From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based models: two training-time strategies (Flow-DPO, Flow-RWR) and one inference-time technique (Flow-NRG). Our experiments show that Flow-DPO outperforms other methods when $\beta$ is fixed. Additionally, Flow-NRG, using a lightweight reward model for noisy videos, enables users to apply custom weightings to multiple alignment objectives during inference.
2 Related Works
Evaluation and Reward Models.
Evaluation models and reward models play a pivotal role in aligning generative models with human preferences. Earlier approaches and benchmarks (Huang et al., 2023; Liu et al., 2024c; Huang et al., 2024) relied on metrics like FID (Heusel et al., 2017) and CLIP scores (Radford et al., 2021) to assess visual quality and semantic consistency. Recent works (Wu et al., 2023a; Xu et al., 2024b; Kirstain et al., 2023; Zhang et al., 2024b; Liang et al., 2024a) have shifted towards utilizing human preference datasets to train CLIP-based models, enabling them to predict preference scores with improved accuracy. With the advent of large vision-language models (VLMs) (Achiam et al., 2023; Wang et al., 2024a), their powerful capabilities in visual understanding and text-visual alignment make them a natural proxy for reward modeling. A common approach involves replacing the token classification head of VLMs with a regression head that predicts multi-dimensional scores for diverse evaluation tasks.
Two main learning paradigms have emerged based on the type of human annotation. The first paradigm relies on point-wise regression, where the model learns to fit annotated scores (He et al., 2024) or labels (Xu et al., 2024a) directly. Another paradigm focuses on pair-wise comparisons, leveraging Bradley-Terry (BT) (Bradley & Terry, 1952) loss to model relative preferences, which is largely unexplored for video reward model. Beyond these methods, some works (Wang et al., 2024b) also leverage the intrinsic reasoning capabilities of VLMs through VLM-as-a-judge (Li et al., 2023; Lin et al., 2024b), where VLMs are adopted to generate preference judgments or scores in textual format through instruction tuning. Despite these promising advances, most existing video reward models primarily focus on legacy video generation models, typically from the pre-Sora (OpenAI, 2024) era, which are constrained by short video durations and relatively low quality. Furthermore, the design and technical choices underlying the vision reward models remain underexplored. Our work seeks to address these limitations by focusing more on modern video generation models and investigating a broader range of reward modeling strategies.
Alignment for Image & Video Generation.
In large language models, Reinforcement Learning from Human Feedback (RLHF) improves alignment with human preferences, enhancing response quality (Ouyang et al., 2022; Jaech et al., 2024). Similar methods have been applied to image generation, using reward models or human preference data to align pretrained models. Key approaches include: (1) direct backpropagation with reward signals (Prabhudesai et al., 2023; Clark et al., 2023; Xu et al., 2024b; Prabhudesai et al., 2024); (2) Reward-Weighted Regression (RWR) (Peng et al., 2019; Lee et al., 2023; Furuta et al., 2024); (3) DPO-based policy optimization (Rafailov et al., 2024; Wallace et al., 2024; Dong et al., 2023; Yang et al., 2024a; Liang et al., 2024b; Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Furuta et al., 2024); (4) PPO-based policy gradients (Schulman et al., 2017) (Black et al., 2023; Fan et al., 2024); and (5) training-free alignment (Yeh et al., 2024; Tang et al., 2024; Song et al., 2023). These methods have successfully aligned image generation models with human preferences, improving aesthetics and semantic consistency. Our work applies the DPO algorithm (Rafailov et al., 2024; Wallace et al., 2024) to flow matching in video generation. Concurrent work (Domingo-Enrich et al., 2024) also explores similar things in image generation. However, they reports worse performance for the DPO-aligned model compared to the unaligned one. We argue that the originally derived Flow-DPO algorithm imposes a stronger KL constraint at lower noise steps, resulting in suboptimal performance. In contrast, using a constant KL constraint significantly improves performance on certain tasks.
There has also been work on aligning video generation models using feedback (Furuta et al., 2024; Li et al., 2024). However, these approaches rely on image reward models (Kirstain et al., 2023; Wu et al., 2023a). Concurrent works (Yuan et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) focus on aligning diffusion-based video generation models by annotating videos generated from earlier open-source models. Our work differs by exploring alignment techniques for advanced flow-based video generation.
@@ 3.1 Human Preference Data Collection
Table 1: Statistics of the collected training dataset. We utilized 12 text-to-video models to generate a total of 108k videos from 16k unique prompts. This process ultimately resulted in 182k annotated triplets, each consisting of a prompt paired with two videos and corresponding preference annotations.
@@
Pre-Sora-Era Models Gen2 (Runway, 2023) 23.06 6k 13k 768 $\times$ 1408 4s
SVD (Blattmann et al., 2023) 23.11 6k 13k 576 $\times$ 1024 4s
Pika 1.0 (Labs, 2023) 23.12 6k 13k 720 $\times$ 1280 3s
Vega (VegaAI, 2023) 23.12 6k 13k 576 $\times$ 1024 4s
Pixverse v1 (PixVerse, 2024) 24.01 6k 13k 768 $\times$ 1408 4s
HiDream (HidreamAI, 2024) 24.01 0.3k 0.3k 768 $\times$ 1344 5s
Modern Models Dreamina (Capcut, 2024) 24.03 16k 68k 720 $\times$ 1280 6s
Luma (LumaLabs, 2024) 24.06 16k 57k 752 $\times$ 1360 5s
Gen3 (Runway, 2024) 24.06 16k 55k 768 $\times$ 1280 5s
Kling 1.0 (Kuaishou, 2024) 24.06 6k 33k 384 $\times$ 672 5s
Pixverse v2 (PixVerse, 2024) 24.07 16k 58k 576 $\times$ 1024 5s
Kling 1.5 (Kuaishou, 2024) 24.09 7k 28k 704 $\times$ 1280 5s
Figure 2: Statistics of our training data.
Existing human preference datasets for video generation (Liu et al., 2024c; Huang et al., 2024; He et al., 2024; Wang et al., 2024b; Xu et al., 2024a) primarily consist of videos generated by earlier open-source models, which are often characterized by low quality, noticeable artifacts, and short video durations. As advancements in video diffusion models (VDMs) continue to elevate the state-of-the-art, the gap between current preference datasets and the capabilities of modern VDMs has become increasingly pronounced. To bridge this gap, we focus on developing a reward model tailored for the latest VDMs. Inspired by Zhang et al. (2024b), we collect a diverse set of prompts from the internet, organizing them into 8 meta-categories (animal, architecture, food, people, plants, scenes, vehicles, objects). These prompts are further expanded using GPT-4o (Achiam et al., 2023). A filtering process is then applied to exclude unsuitable prompts (e.g., repetitive, irrelevant to motion, unsafe content). Subsequently, we employ our internal prompt rewriter to refine and enhance the collected prompts, resulting in a final dataset of 16k detailed prompts. Leveraging these prompts, we select 12 text-to-video generation models with various capabilities, producing 108k videos and constructing 182k triplets, each comprising a prompt and two corresponding videos generated by different video generation models. Comprehensive statistics of the dataset are provided in Tab. 1 and Fig 2.
To annotate these text-video triplets, we hire annotators to perform pairwise assessments across three critical dimensions: Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA). Details are available in appendix C. For each dimension, annotators are presented with a prompt and two generated videos, and asked to provide choices indicating their preference (A wins/Ties/B wins). Each sample is independently assigned to three annotators to ensure robustness. In cases of disagreement, an additional annotator will be engaged to resolve conflicts. This effort results in a dataset of 182k high-quality, multi-dimensional preference annotations, which were utilized for training the reward model. In addition to pairwise annotations, we also conduct pointwise annotations using a similar procedure, where annotators rate each video on a Likert scale from 1 to 5 for the same three dimensions (i.e., VQ, MQ, TA). This dual annotation setup enables us to explore the advantages and limitations of pairwise preferences versus pointwise scores as annotation strategies.
In addition to the training data, we construct a validation set comprising 13k annotated triplets, following a similar annotation protocol. The prompts in the validation set are strictly excluded from the training set to ensure generalization. The validation set serves as a reliable benchmark to monitor the convergence during training.
3.2 Reward Model Learning
Following prior works that leverage Vision-Laguage Models (VLMs) for related tasks (He et al., 2024; Wang et al., 2024b; Xu et al., 2024a), we adopt Qwen2-VL-2B (Wang et al., 2024a) as the base model for our video reward framework. While existing studies have demonstrated the effective application of reward models in both evaluation (Wu et al., 2023b, a; He et al., 2024) and optimization (Lee et al., 2023; Wallace et al., 2024; Xu et al., 2024b; Prabhudesai et al., 2024), the underlying design choices remain insufficiently explored. We start with a Bradley-Terry style reward model as our baseline, exploring how technical decisions affect the video reward model’s performance.
@@ 3.2 Reward Model Learning
We first explore the merits of adopting a Bradley-Terry pairwise comparison approach compared to a pointwise score regression approach for training reward models. The BT model (Bradley & Terry, 1952) uses a pairwise log-likelihood loss to capture the probability that one video is preferred over another, enabling the prediction of rewards based on the relative preferences. Conversely, the pointwise regression model directly estimates the absolute quality scores by minimizing the MSE loss. The loss functions for these methodologies are formally defined as follows:
$\mathcal{L}_{BT}=-\mathbb{E}_{({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})\sim\mathcal{D}}\left[\log\left(\sigma\left(r({\bm{x}}^{w}_{0},{\bm{y}})-r({\bm{x}}^{l}_{0},{\bm{y}})\right)\right)\right]$ (1)
$\mathcal{L}_{reg}=\mathbb{E}_{({\bm{y}},{\bm{x}}_{0},z)\sim\mathcal{D}}\left[\|r({\bm{x}}_{0},{\bm{y}})-z\|^{2}\right]$ (2)
In these formulations, $({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})$ denotes pairwise annotations, where ${\bm{y}}$ is the input prompt and ${\bm{x}}^{w}_{0}$ and ${\bm{x}}^{l}_{0}$ are the two generated videos with corresponding preference annotations. Meanwhile, $({\bm{y}},{\bm{x}}_{0},z)$ represents pointwise annotations, with $z$ being the assigned quality score. The function $r$ signifies the optimized reward function.
Given that our training dataset includes both pointwise scores and pairwise preferences, annotated by the same group of annotators on the same dataset, this setup enables a direct comparison between the two paradigms. To conduct this comparison, we train both types of reward models on the training data with varying proportions of annotated data. Their performance is assessed by evaluating the average accuracy of the best-performing checkpoints on the validation set. Notably, for the BT reward model, we exclude samples labeled as ”Ties” and only utilize samples annotated as ”A Wins” or ”B Wins” since the BT framework is not well-suited to handle tied data. The comparison results are illustrated in Fig. 3.
Figure 3: Accuracy comparison between the BT and regression reward models across varying training data fractions (log scale).
We observe that as the dataset size increases, both the Bradley-Terry reward model and the regression reward model exhibit notable improvements in accuracy, with the BT model consistently outperforming the regression model. This disparity can be attributed to the inherent characteristics of pointwise and pairwise annotations. Pairwise annotations are inherently more effective at capturing relative differences between samples. For instance, even when two videos receive the same score in pointwise annotations, annotators can still discern their relative quality in pairwise comparisons. This ability grants the BT reward model a distinct advantage when the dataset size is limited. As the dataset expands, the performance gap between the two models narrows, as the regression model benefits from the increased data to learn more accurate ranking relationships. Additionally, the annotation process itself varies in complexity between the two approaches. Pairwise preference labels are generally easier for annotators to agree upon compared to assigning precise scores in pointwise annotations. Recent studies (Wang et al., 2024d, c) have demonstrated that with carefully curated pointwise annotations, regression models can achieve performance levels comparable to BT reward models. We posit that as both the quantity and quality of the data improve, the performance differences between these approaches become negligible. While in the context of our current training data, the BT reward model emerges as the better-performing choice.
@@
While the BT model is commonly used to capture human preferences from chosen-rejected pairs, the importance of tie annotations is often overlooked. Inspired by recent works (Liu et al., 2024a), we use the Bradley-Terry model with ties (BTT) (Rao & Kupper, 1967), an extension of BT that accounts for tied preferences. For a given prompt ${\bm{y}}$ and two generated videos ${\bm{x}}^{1}_{0}$ and ${\bm{x}}^{2}_{0}$, the BTT model defines the probabilities of each possible preference as follows:
$P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{(\theta^{2}-1)e^{r({\bm{x}}_{0}^{A},{\bm{y}})}e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\left(e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)\left(\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}\right)}$ (3)
$P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}+\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}}$ (4)
$P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})$ $=\frac{e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}{\theta e^{r({\bm{x}}_{0}^{A},{\bm{y}})}+e^{r({\bm{x}}_{0}^{B},{\bm{y}})}}$ (5)

### v2

@@ Abstract
Video generation has achieved significant advances through rectified flow techniques, but issues like unsmooth motion and misalignment between videos and prompts persist. In this work, we develop a systematic pipeline that harnesses human feedback to mitigate these problems and refine the video generation model. Specifically, we begin by constructing a large-scale human preference dataset focused on modern video generation models, incorporating pairwise annotations across multi-dimensions. We then introduce VideoReward, a multi-dimensional video reward model, and examine how annotations and various design choices impact its rewarding efficacy. From a unified reinforcement learning perspective aimed at maximizing reward with KL regularization, we introduce three alignment algorithms for flow-based models. These include two training-time strategies: direct preference optimization for flow (Flow-DPO) and reward weighted regression for flow (Flow-RWR), and an inference-time technique, Flow-NRG, which applies reward guidance directly to noisy videos. Experimental results indicate that VideoReward significantly outperforms existing reward models, and Flow-DPO demonstrates superior performance compared to both Flow-RWR and supervised fine-tuning methods. Additionally, Flow-NRG lets users assign custom weights to multiple objectives during inference, meeting personalized video quality needs.
@@ 1 Introduction
Advancements in video generation have led to powerful models [57, 33, 32, 5] that produce realistic details and coherent motion. Despite this, current systems still face challenges like unstable motion, imperfect text-video alignment and insufficient alignment with human preferences [84]. In language model and image generation, reinforcement learning from human feedback (RLHF) [52, 90, 71] has proven effective in improving response quality and aligning models with user expectations.
However, applying RLHF to video generation is still remains in its infancy. A major obstacle is the lack of a reliable reward signal. Existing preference datasets [19, 73, 47, 76] were collected on earlier T2V models that produced short, low-resolution clips. Reward models trained on such data may miss fine spatial detail and long-range dynamics, while over-penalising glitches that current T2V models already suppress. In addition, the design space of VLM-based reward models remains under-explored, leading to sub-optimal annotation paradigms, reward hacking issues, and entangled multi-attribute scores. The resulting supervision is therefore noisy, biased, and easily exploited during RLHF.
@@
A second challenge arises from the internal mechanisms of cutting-edge video generation models. Many modern systems employ rectified flow [48, 45], predicting velocity rather than noise. Recent studies [73, 85, 47, 76] have tested DPO [61, 71] and RWR [54, 36, 17] on diffusion-based video generation approaches. However, adapting existing alignment methods to flow-based models introduces new questions. A recent attempts [13] for flow matching based DPO even degrade quality compared with the unaligned baseline.
@@
Figure 1: Overview of Video Alignment Framework. (a) Human Preference Annotation. We construct 182k prompt-video triplets, each annotated on Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA). (b) Reward Modeling. A VLM-based reward model is trained under the Bradley-Terry-Model-with-Ties formulation. (c) Video Alignment. From a unified RL perspective, we introduce three alignment algorithms for flow-based video generation: Flow-DPO, Flow-RWR, and Reward Guidance (Flow-NRG), and provide a systematic comparison.
To address these challenges, we present a comprehensive investigation into aligning advanced flow-based video generation models with human preferences, as shown in Fig. 1. We first collect 16k high-quality prompts and render them with 12 representative T2V models, producing 182k annotated examples across three key dimensions: Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA). We then develop a multi-dimensional video reward model, systematically analyzing how different annotations and design choices affect its performance. These dimensions can be aggregated into a total reward that reflects the overall preferences of humans.
Armed with a reliable reward model, we revisit RLHF algorithms for rectified flow. From a unified reinforcement learning perspective that maximizes reward with KL regularization, we derive two training-time strategies—Flow-DPO and Flow-RWR, and an inference-time technique, Flow-NRG. We discover that a simple extension of Diffusion-DPO performs poorly because its timestep-dependent KL term pushes the policy to overfit the objective at higher noise levels, making the model vulnerable to reward hacking. In contrast, our Flow-DPO removes this time-dependent term and keeps the KL weight constant, showing robust performance across all dimensions.
Flow-NRG is an efficient inference-time alignment algorithm that applies reward guidance directly to noisy videos during the denoising process. It allows users to apply arbitrary weightings to multiple alignment objectives during inference, eliminating the need for extensive retraining to meet personalized user requirements.
@@
  • Large-Scale Preference Dataset: We create a 182k-sized, multi-dimensional, human-labeled video generation preference dataset from 12 modern video generation models.
@@
  • VideoGen-RewardBench: We create a benchmark for modern reward models by annotating prompt-video pairs from VideoGen-Eval. This dataset, consisting of diverse prompts and videos generated by modern VDMs, results in 26.5k annotated video pairs with preference labels.
@@
  • Flow-Based Alignment: From a unified RL perspective, we introduce two training-time alignment strategies (Flow-DPO, Flow-RWR) and one inference-time technique (Flow-NRG). Experiments show that Flow-DPO with a fixed KL term outperforms other methods. And Flow-NRG allows custom weightings of multiple alignment objectives during inference.
2 Related Work
Reward modeling [75, 77, 31, 86, 41] trains CLIP-based models on human preference datasets, while newer approaches use VLMs with regression heads to predict multi-dimensional scores. Learning paradigms include point-wise regression [19, 76] and pair-wise comparison via Bradley-Terry loss. However, most video reward models focus on short, low-quality videos from the pre-Sora era [51] and lack rigorous evaluation of design choices. We address these limitations by targeting modern video generation and exploring broader reward modeling strategies. For alignment, image generation has adopted RLHF-style techniques such as reward backpropagation [58, 77], RWR [36], DPO [61, 71], PPO [2, 16], and training-free methods [80]. Concurrent efforts [17, 47, 76, 85, 73] extend DPO/RWR to diffusion-based video models using old generation models or image-level rewards. We build on this by extending DPO to flow-based video generation, proposing Flow-DPO. More comprehensive discussions of related work can be found in Appendix A.
@@ 3 VideoReward
Robust RLHF begins with a reward model that faithfully mirrors human preferences, yet existing efforts are limited in two key respects: (1) Data: existing video-preference datasets were curated for earlier T2V models and mismatch with what users prefer in modern video generation models; (2) Model design: the key technique choices for VLM-based reward models remain largely uncharted. Fig. 1 (a), (b) summarises our solution. We first build a large-scale preference dataset collected with state-of-the-art T2V models; then perform a systematic study of reward-modeling design.
@@ 3.1 Human Preference Data Collection
Table 1: Statistics of the collected training dataset. We utilize 12 T2V models to generate 108k videos from 16k unique prompts, resulting in 182k annotated triplets. Each triplet consists of a prompt paired with two videos, and corresponding preference annotations.
@@
Pre-Sora-Era Models Gen2 [63] 23.06 6k 13k 768 $\times$ 1408 4s
SVD [3] 23.11 6k 13k 576 $\times$ 1024 4s
Pika 1.0 [34] 23.12 6k 13k 720 $\times$ 1280 3s
Vega [70] 23.12 6k 13k 576 $\times$ 1024 4s
PixVerse v1 [56] 24.01 6k 13k 768 $\times$ 1408 4s
HiDream [21] 24.01 0.3k 0.3k 768 $\times$ 1344 5s
Modern Models Dreamina [6] 24.03 16k 68k 720 $\times$ 1280 6s
Luma [50] 24.06 16k 57k 752 $\times$ 1360 5s
Gen3 [64] 24.06 16k 55k 768 $\times$ 1280 5s
Kling 1.0 [33] 24.06 6k 33k 384 $\times$ 672 5s
PixVerse v2 [56] 24.07 16k 58k 576 $\times$ 1024 5s
Kling 1.5 [33] 24.09 7k 28k 704 $\times$ 1280 5s
Existing human preference datasets for video generation [49, 26, 19, 73, 76] were primarily built on early, low-resolution T2V models that produced short, artifact-laden clips. As VDMs continue to evolve, modern T2V models, however, generate longer, higher-fidelity videos with smoother motion. Consequently, legacy datasets no longer accurately reflect what users prefer today. Reward models trained on such collections may miss fine spatial details and long-range dynamics while over-weighting temporal glitches already mitigated by current models. To bridge this gap, we develop a new preference dataset expressly for state-of-the-art VDMs.
Prompt Collection and Video Generation.
We collect diverse prompts from the Internet, categorize them into 8 meta-categories—animal, architecture, food, people, plants, scenes, vehicles, objects—and expand them with GPT-4o. After removing repetitive, irrelevant, or unsafe entries, we refine them with our in-house prompt rewriter, yielding 16000 high-quality prompts. 12 T2V models of varying capabilities then render these prompts into 108k videos, which we pair to form 182k triplets, each comprising a prompt and two corresponding videos from distinct VDMs. Comprehensive dataset statistics are provided in Tab. 1 and Fig. 8.
Multi-dimensional Annotation.
Professional annotators were hired to view each triplet and record pairwise preferences(A wins / Ties / B wins) separately for Visual Quality (VQ), Motion Quality (MQ), and Text Alignment (TA), producing a three-label vector per triplet. The same annotators also assign 1–5 Likert scores to each individual video, enabling later studies that compare pointwise and pairwise supervision. We reserve 13 000 triplets whose prompts never appear in training as a validation set. Detailed annotation protocols are provided in Appendix G.
3.2 Reward Modeling
Prior works [19, 73, 76] on VLM-based reward models have proved effective for both evaluation [75, 19] and optimization [36, 71, 59]. However, their core design choices remain insufficiently explored, and current methods still suffer from issues like annotation paradigms, reward hacking and entangled multi-attribute signals. We select the lightweight Qwen2-VL-2B [72] as our backbone, and systematically examine three key designs and demonstrate how each yield a cleaner, more reliable reward signal for RLHF. We also conduct a comprehensive ablation study on the evaluation benchmark in Table 6 of Appendix E to further validate the key design choices.
@@ 3.2 Reward Modeling
We first investigate two reward learning paradigms: the Bradley-Terry (BT) model [4] and pointwise score regression. The BT model formulates preference learning as a probabilistic ranking task. Given a prompt $\bm{y}$} and paired videos $({\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})$, it optimizes $\mathcal{L}_{BT}=-\mathbb{E}\left[\log\left(\sigma\left(r({\bm{x}}^{w}_{0},{\bm{y}})-r({\bm{x}}^{l}_{0},{\bm{y}})\right)\right)\right]$, where the expectation is taken over $({\bm{y}},{\bm{x}}^{w}_{0},{\bm{x}}^{l}_{0})\sim\mathcal{D}$}.
In contrast, score regression directly predicts a scalar quality score $z\in\mathbb{R}$ using the MSE loss: $\mathcal{L}_{reg}=\mathbb{E}\left[\|r({\bm{x}}_{0},{\bm{y}})-z\|^{2}\right],$ where the expectation is taken over $({\bm{y}},{\bm{x}}_{0},z)\sim\mathcal{D}$}.
Since our training dataset includes both pointwise scores and pairwise preferences from the same annotators, we can directly compare between the two annotation paradigms. We train both types of reward models on increasing subsets of the training set and report the best validation accuracy averaged over VQ, MQ, and TA. Fig. 3 presents these results.
Figure 2: BT vs. Regression. Accuracy curves across log-scaled data fractions.
Figure 3: BT vs. BTT. Score-difference distributions ($\Delta r$) for BT (left) and BTT (right).
As the dataset grows, both the BT and regression-style models improve in accuracy, while the BT model remains consistently superior. This advantage stems from the nature of pairwise annotations, which capture subtle relative distinctions more effectively. Even when two videos receive identical pointwise scores, annotators can still differentiate subtle quality differences.
@@
While vanilla BT model is widely used to capture human preferences from chosen-rejected pairs, the importance of tie annotations is often overlooked. Inspired by recent work [46], we adopt Bradley-Terry model with ties (BTT) [62], an extension that accounts for tied preferences. Formally, BTT defines a tripartite preference distribution:
$P_{\theta}(c|{\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})=\begin{cases}\displaystyle\frac{(\theta^{2}-1)\,\exp({r_{A}})\exp({r_{B}})}{(\exp({r_{A}})+\theta\exp({r_{B}}))\,(\theta\exp({r_{A}})+\exp({r_{B}}))}&\text{Tie }(x^{A}_{0}=x^{B}_{0}),\\[8.0pt] \displaystyle\frac{\exp({r_{A}})}{\exp({r_{B}})+\theta\exp({r_{A}})}&\text{A preferred }(x^{A}_{0}\succ x^{B}_{0}),\\[8.0pt] \displaystyle\frac{\exp({r_{B}})}{\theta\exp({r_{A}})+\exp({r_{B}})}&\text{B preferred }(x^{B}_{0}\succ x^{A}_{0}).\end{cases}$ (1)